OpenStudy (anonymous):

Verify the identity. cotangent of x to the second power divided by quantity cosecant of x plus one equals quantity one minus sine of x divided by sine of x

OpenStudy (anonymous):

\[\cot^2x/\csc x +1 = 1 - \sin x/\sin x\]

OpenStudy (welshfella):

try changing LHS to sine and cosines

OpenStudy (anonymous):

(cos^2(x)/sin^2(x))/1/sin(x) = (cos^2(x)/sin^2(x))/sin(x)

OpenStudy (welshfella):

just checking is the question